A social psychologist wants to test if extroversion is significantly related to a person’s general health. He gets extroversion scores and a measure of general health from the participants’ medical providers. The data are listed below. Higher health extroversion scores mean greater extroversion, while high health scores = good health. Complete the appropriate statistical test and hypothesis testing using the 4 step approach. State your conclusion in APA format. Alpha level = .05. Participant extrov X health Y X squared Y squared (X) (Y) A 101 33 3333 B 93 40 3720 C 85 24 2040 D 111 46 5106 E 104 30 3120 F 89 28 2492 Total
Answer #1

Answer:

Given Data

solution 8 = n=6 Ex-583 Ey= = 201 Exy = 19311 sst = 282 ssy = 310 sexy = {(x1 - x) lyi-y) = {xy-nxy 97.1667 ។ Ey 201 = 33.5 STest statistico -tn-2 1-82 1.8972 = 5.9992 0-94866-2 ſi-0-94852 0.3164 p-value=0.002 tcritical=tn-2 , 2/2 = t4, 0.025 = 2-770****Please like it..

It is important to me....

Thank you for supporting me....